OpenStudy (anonymous):

What are the foci of the ellipse given by the equation 4x^(2)-32x+9y^(2)+18y=37?

OpenStudy (anonymous):

Yay i love these equations....not alright so you have to solve the equation first? do you know how to do that

OpenStudy (anonymous):

no

OpenStudy (anonymous):

alright do you know the equation for an ellipse

OpenStudy (dumbcow):

heres a reference i've made for these types of equations

OpenStudy (anonymous):

first what you want to do is try to create the equation of an ellipse. \[(x-h)^2/a^2 + (y-k)^2/b^2\]

OpenStudy (anonymous):

=1

OpenStudy (dumbcow):

you have to do the same steps i did for prev problem

OpenStudy (anonymous):

Thats what I am stuck on, idk how to put it into a equation and dumbcow i didn't understand your way

OpenStudy (anonymous):

so first you have your numbers grouped but you can divide out a common factor of 4 from x and 9 from y getting \[4(x^2-8x)+9(y^2+2y)=37\]

OpenStudy (dumbcow):

theres really no other way to do it though to get it in standard form

OpenStudy (anonymous):

now in order to get x-h^2 and (y-k)^2 you can obviously see that you'll need to complete the square inorder to factor it

OpenStudy (anonymous):

I forget how to complete the square

OpenStudy (anonymous):

sooooo \[4(x^2-8x+16)+9(y^2+2y+1)=37\] however whatever you add to one side you must add to the other and you have to take account for the factored out integers so 4*16 and 9*1 you'll get \[4(x-8x+16)+9(y^2+2x+1)= 37+(16*4)+(9*1) \]

OpenStudy (dumbcow):

bandgeek are you sure its not supposed to be equal to -37

OpenStudy (anonymous):

yeah sorry, it equals -37

OpenStudy (anonymous):

when completing the square you take your base x or your second term (i'm not 100% what they call it when they teach this) and you divide by 2 and square the number

OpenStudy (anonymous):

so for the first one : -8x/2 = -4^2 = 16

OpenStudy (anonymous):

now when you factor this it will be a perfect square and you can now factor it to (x-4)^2

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

let m know when you get back or what not?